Circular No. 8882

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2007 J8–J12 AND C/2007 K7–K18 (SOHO) Additional near-sun comets (cf. IAUC 8853) have been found on SOHO

website images — all being Kreutz sungrazers except for C/2007 K12 (Meyer group). K. Battams reports that C/2007 J8, K7, K9, and K11 were slightly diffuse and very faint (mag 8) with no tail. C/2007 J9 was stellar (mag ∼ 6.5) in C3 images, and very slightly diffuse with a hint of tail in C2 images. C/2007 J10, K13, K15, and K16 were slightly diffuse and faint (mag ∼ 7.5). C/2007 J11 was stellar in C3 images, and was of mag ∼ 6.5 with a hint of tail in C2 images. C/2007 J12 was extremely faint in C3 images and “very faint/diffuse” (mag 8) in C2 images. C/2007 K8 was stellar in appearance in C3 images, and slightly diffuse (mag ∼ 7) with a hint of tail in C2 images. C/2007 K10 was stellar in appearance and faint in C3 images, and slightly diffuse and elongated (mag ∼ 7.5) in C2 images. C/2007 K12 was stellar in appearance, elongated (but not in the direction of motion), and of mag ∼ 6.5. C/2007 K14 was extremely faint (mag ∼ 8.5) and fairly diffuse. C/2007 K17 was very faint (mag ∼ 8), elongated, and diffuse. C/2007 K18 was faint and stellar in appearance in C3 images, and slightly diffuse (mag ∼ 7) with a hint of tail in C2 images.

Comet  2007 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2007 J8  May 12.913 3^{h}20.^{m}9 +16^{o}49′ C2  BZ 2007-T56
C/2007 J9  13.904 3 28.4 +16 14 C3/2 SY 2007-T56
C/2007 J10 14.135 3 25.7 +17 01 C2 BZ 2007-T56
C/2007 J11  15.433 3 40.0 +15 36 C3/2 BZ 2007-T56
C/2007 J12  15.679 3 38.4 +16 02 C3/2 BZ 2007-T56
C/2007 K7 16.854 3 34.3 +17 31 C2 BZ 2007-T57
C/2007 K8  17.196 3 40.9 +16 47 C3/2 BZ 2007-T57
C/2007 K9 17.646 3 38.5 +17 46 C2 BZ 2007-T57
C/2007 K10  17.779 3 43.4 +16 55 C3/2 BZ 2007-T77
C/2007 K11  18.271 3 41.1 +17 53 C2  BZ 2007-T77
C/2007 K12  18.396 3 41.9 +20 08 C2  HS 2007-T77
C/2007 K13  21.175 3 52.4 +18 30 C2  TH 2007-T77
C/2007 K14  21.183 3 52.5 +18 29 C2  BZ 2007-T78
C/2007 K15  22.200 3 54.8 +18 34 C2  RK 2007-T78
C/2007 K16  22.292 3 55.9 +18 38 C2  HS 2007-T78
C/2007 K17  24.635 4 05.2 +19 07 C2  TH 2007-T78
C/2007 K18  26.263 4 11.9 +19 06 C3/2 MM 2007-T78
2007 October 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green